An electric flux of 143 N.m2/C passes through a flat horizontal surface that has an area of 0.62 m2. The flux is due to a uniform electric field. What is the magnitude of the electric field if the field points 15o above the horizontal?
Answer #1

φ = 143 N.m2/C, A = 0.62 m2, θ = 15o

find E

φ = EAsinθ, ∴E = φ/(Asinθ) = 891 N/C